OpenStudy (anonymous):

8x/7 + 9 = 30. cant figure answer out

OpenStudy (anonymous):

8x/7 + 9 = 30 First subtract 9 from both sides. You get 8x/7 = 21 Then, multiply both sides by 7. You get 8x = 147 Then divide by 8 on both sides. You get x = 147/8